If Z is a standard normal random variable, what is the value z for which P(0 ≤ Z ≤ z) = 0.35


1
Expert's answer
2021-01-11T15:35:46-0500

We have that "P(0 \u2264 Z \u2264 z) = 0.35"

"P(0 ≤ Z ≤ z) = P(Z\le z) - P(Z\le 0)=P(Z\le z)-0.5 = 0.35"

Hence

"P(Z\le z)= 0.35+0.5=0.85 \implies z = 1.3" from normal table we find corresponding z value


Answer: 1.3